# Engagement Maximization

### v5

@@ 1 Introduction
Free-to-use online platforms such as Facebook, Instagram, Youtube, and Pinterest are used by billions of people worldwide and earn substantial profits by displaying advertisements to their users. Their business models are powered by personalized recommendation algorithms that seek to maximize the “engagement” of each user by selectively displaying content (Lada et al. (2021), Sequoia (2018)). Large quantities of computing and other resources have been invested by these firms to develop algorithms that can predict billions of users’ preferences in real time.
@@
A second and more subtle challenge is to manage the algorithm’s own “incentives”. Platforms often design these algorithms with the goal of providing reliable and wide-ranging content. However, subsequent learning based on user interactions can lead algorithms to maximize engagement by leading users down niche “rabbit holes ” (WSJ (2021)). This can lead to short term engagement gains but in the long-run will cause users to question the credibility of the platform. It is not obvious ex-ante whether concerns about credibility meaningfully constrain the algorithm design.
In this paper, we study the optimal design of sequential information presentation from a principal-agent perspective. The principal (the platform) provides the agent (the user) with information. The agent derives value from the information, and chooses when to stop engaging with the platform. [^1] We show that it is without loss of generality to assume the agent does not process the information selectively; that is, the agents attends to the information the principal provides in equilibrium. The rate at which the agent can process information is limited, and the agent experiences an opportunity cost of time spent on the platform net of any utility from using the platform. The principal’s goal is to maximize the attention the agent allocates to the platform (which under some conditions will be equivalent in equilibrium to maximizing the stopping time), and the principal accomplishes this by choosing the nature of the information provided to the agent. The key modeling assumptions we impose are that the principal knows perfectly the agent’s preferences and that the principal can flexibly manipulate the entire information flow. These assumptions capture the state platforms aim to reach with unlimited data and computing power.
@@
    Leading Case: Revenue from all engagement. Suppose that the principal’s engagement measure is identical to the informativeness measure. We interpret this case as one in which the agent is exposed to ads in proportion to the context she receives, with no measurement of whether she attends to the ads. In this case, the optimal information structure is narrow but deep: the principal only provides information about the agent’s interested dimensions. This information is provided in a rare but chunky way, leading the agent to overspend their time and overdose information relative to the agent-optimal benchmark. This result rationalizes the commonly criticized rabbit-holing and polarization behavior of the recommendation algorithms when they target the maximization of engagement metrics.
    Leading Case: Revenue from decision-irrelevant engagement. Now suppose that the principal’s revenue comes from engagement with information that is irrelevant for decision. We interpret this case as one in which the agent is exposed to decision-irrelevant ads, and the principal benefits only when the agent engages with these ads (e.g. if ad revenue comes from the click of an ad.). In this case, we show that the optimal information structure is broad but noisy: the principal provides excess information (e.g. sponsored content) that the agent views as pure noise, intermixed with just enough decision-relevant information to keep the agent engaged. This result rationalizes the click-baiting and stealth marketing behavior of the recommendation algorithms when they target the maximization of clicks.
@@
    The intuition for the result is simple: We show that the principal-optimal policy leaves the agent with no surplus at the prior. If the agent were to become “more certain than” the prior at an interim stage, the surplus from future information would be negative, and the agent would stop engaging with the principal. On the other hand, if there agent were to become “less certain than” the prior at an interim stage, her surplus from future information would be positive. At this point, the principal would be tempted to have the algorithm adjust and extract more surplus, at the agents’ expense. Thus, the only possible interim beliefs are those that are exactly “as certain as” the prior belief. Our main result in Section 4 is a formalization of this idea, along with a precise definition of what it means to be more certain, less certain, or as certain as the prior.
After presenting our main results , we consider several extensions of our baseline model. First, we consider the possibility that the agent does not fully attend to the signals sent by the principal. We show that it is without loss of generality to assume the agent pays full attention to the signals provided. Second, we show that the usual Bayesian persuasion assumption of commitment to a signal structure is without loss in our setting. Third, we consider the case in which there is an increasing opportunity cost of time and/or decreasing marginal utility from using the platform. This case is less tractable; to make progress, we impose additional symmetry assumptions, and derive conditions under which our main results continue to hold. Fourth, we discuss the importance of the role of jumps in beliefs in our model. The optimal belief process in our baseline model will involve rare but informative signal realizations that cause beliefs to jump. This result speaks to the sensational nature of the information that the principal optimally provides the agent. To clarify this point, we extend our analysis to the case in which the size of jumps in beliefs is arbitrarily restricted. We characterize how the restriction on jump size restricts the set of implementable strategies, and show that under certain circumstances, if beliefs must follow a continuous process, the principal and agent are in fact fully aligned and it is instead the agent who receives the full surplus. Fifth, we consider a modification of the model in which the user has unlimited information processing capacity and the platform maximizes time spent (as opposed to attention). We show that some of our results continue to apply to this modified model.
Throughout the paper, we discuss the example an ad-supported internet platform engaging a user. However, we pair this example with a second example, that of a teacher who seeks to maximize the engagement of a student who cares only about passing a test. [^2] We are grateful to Emir Kamenica for suggesting this alternative setting as an application of our model. Our theoretical framework applies to this setting without modification, and we discuss both examples to emphasize that engagement maximization is relevant in a wide range of contexts. As an example, the distinction we make between per-impression and per-click advertising in the context of an internet platform maps, in the teacher-student context, to the degree to which the teacher values test-relevant vs. test-irrelevant information. A teacher who values student learning but not test performance will provide just enough test-specific information to engage a student who cares only about the test, while providing additional content that is not part of the test. In contrast, a teacher who values all learning, regardless of whether it is covered on a test, will optimally provide only test-relevant information, while forcing students to learn more test-relevant information than they would choose to learn on their own.
@@ 1.1 Related Literature
Our paper contributes to several strands of literature on the dynamic provision of information. Viewing our principal as a media company, our model is related to work on models of media bias (see Gentzkow et al. (2015) for a survey). We share with Kleinberg et al. (2022) an interest in explaining why the users of internet platforms would engage heavily with those platforms while perceiving themselves as gaining little from doing so. We derive this outcome as a result of strategic behavior by rational agents with conflicting incentives; those authors emphasize the time-inconsistency of user preferences. We share with Acemoglu et al. (2021) an emphasis on explaining what kind of information is available on internet platforms; our analysis focuses on content selection algorithms, whereas their analysis focuses on information sharing between users.
@@
Closely related to our work is the literature on dynamic Bayesian persuasion (e.g. Ely (2017); Renault et al. (2017); Ely and Szydlowski (2020); Orlov et al. (2020); Che et al. (2020)) that build on the static model of Kamenica and Gentzkow (2011). Our setting differs from most of these papers in that our principal’s payoff solely depends on the total attention paid by the agent, while the agent obtains pure instrumental value of information from solving a general decision problem. [^3] Our model is equivalent to one where the principal’s only goal is to ensure the agent continues to pay attention, as in Kawamura and Le Quement (2019). However, our principal presents unbiased information, and hence is not engaged in cheap talk (Crawford and Sobel (1982); Cheng and Hsiaw (2022)). Most of these papers focus on the “Bayesian persuasion” settings where the principal’s payoffs directly depend on the agent’s action and the maximization of “engagement” is a side effect. Two exceptions are Knoepfle (2020)and Koh and Sanguanmoo (2022), who also focus on maximizing “engagement” or “attention,” but without an information processing constraint. [^4] Knoepfle (2020)’s main focus is on the competition between multiple senders. The single sender case of Knoepfle (2020) is the closest to ours. In both papers, the lack of an information constraint leads to the optimal persuasion strategy fully revealing the state. Our analysis emphasizes the beliefs the agent will arrive at, and how these differ between the principal’s optimal strategy and an agent-optimal benchmark, a comparison that is not possible absent information processing constraints. We elaborate further on the connection between our model and these models in Section 5.5. We assume both players are long-lived and have commitment power, in contrast to the limited commitment settings in Orlov et al. (2020); Che et al. (2020) or the myopic agent settings in Ely (2017); Renault et al. (2017). Interestingly, unlike the findings in these papers, neither commitment power nor forward-lookingness of the agent is necessary for sustaining our equilibrium strategy (see Section 4.2).
@@
Our model predicts gradual information revelation over time, which is a feature shared by many of the dynamic information design models (e.g. Ely et al. (2015); Hörner and Skrzypacz (2016); Che et al. (2020); Orlov et al. (2020)). However, unlike these papers, the gradual nature of belief evolution in our model arises from the agent’s information processing constraint, as opposed to a desire to maximize suspense or address problems of limited commitment. In particular, Che et al. (2020) predicts that the agent’s belief involves Poisson jumps and a drift, while our optimal strategy admits Poisson jumps without a drift.
@@
Formally, our approach is a principal-agent version of Hébert and Woodford (2023). Those authors consider a model in which a single decision maker chooses both what information to acquire and when to stop and act, whereas in our model the principal chooses the information and the agent chooses when to stop and act. We compare our model to a benchmark in which the agent chooses both the information and when to stop and act; this benchmark is characterized by results found in Hébert and Woodford (2023). We follow Hébert and Woodford (2023) in assuming that the principal can choose any stochastic process for the agent’s beliefs, subject only to the martingale requirement (which is imposed by Bayesian updating) and the upper bound on the agent’s attention. We model this upper bound using a “uniformly posterior-separable” information cost, in the terminology of the rational inattention literature (Caplin et al. (2022)). [^5] Examples of such information costs include mutual information, as applied in Sims (2010), as well as other proposed alternatives (Hébert and Woodford (2021), Bloedel and Zhong (2020)).
@@
The rest of the paper is organized as follows. We begin in section 2 by describing the basic environment of our model. Section 3 characterizes optimal policy in our baseline model, with an emphasis on the beliefs the agent will hold when stopping. Section 4 discusses the dynamics of beliefs and the role of commitment. Section 5 discusses several extensions of our baseline model, and in Section 6 we conclude.
@@ 2.1 The Agent’s Problem
We study the problem of a rational, Bayesian agent receiving signals about an underlying state for the purpose of taking an action. We model information acquisition as a continuous time process, building on results in Hébert and Woodford (2023) and Zhong (2022).
Let $X$ be a finite set of possible states of nature. The state of nature is determined ex-ante, does not change over time, and is not known to the agent. Let $q_{t}\in\mathcal{P}(X)$ denote the agent’s beliefs at time $t\in[0,\infty)$, where $\mathcal{P}(X)$ is the probability simplex defined on $X$. We will represent $q_{t}$ as vector in $\mathbb{R}_{+}^{|X|}$ whose elements sum to one, each of which corresponds to the likelihood of a particular element of $X$, and use the notation $q_{t,x}$ to denote the likelihood under the agent’s beliefs at time $t$ over the true state being $x\in X$.
At each time $t$, the agent can either stop and choose an action from a finite set $A$, or continue to acquire information. Let $\tau$ denote the time at which the agent stops and makes a decision, with $\tau=0$ corresponding to making a decision without acquiring any information. The agent receives utility $u_{a,x}$ if she takes action $a$ and the true state of the world is $x$, and pays a flow cost of delay per unit time, $\bar{\kappa}>0$, until an action is taken. This flow cost captures the opportunity cost of the agent’s time net of whatever utility is gained as part of the information acquisition process (e.g. from using the principal’s platform). In 5.3, we provide an extension where there is an increasing time-dependent flow cost of delay $\kappa(t).$ [^6] Note that we assume that the agent does not discount the utility from acting. Discounting, in our context, has the effect of changing the relative values of acting and never acting, and for this reason would complicate our analysis; see the more extensive discussions in Hébert and Woodford (2023) and Zhong (2022) on this point.
Let $\hat{u}(q^{\prime})$ be the payoff (not including the cost of delay) of taking an optimal action under beliefs $q^{\prime}\in\mathcal{P}(X)$:
$\hat{u}(q^{\prime})=\max_{a\in A}\sum_{x\in X}q^{\prime}_{x}u_{a,x}.$
In what follows, the convex function $\hat{u}:\mathcal{P}(X)\rightarrow\mathbb{R}$ will summarize the value the agent places on information. This function, as opposed to the utility function $u_{a,x}$, can be viewed as the primitive “value of information when acting” in our model; nothing in our analysis will depend on the nature of the action space $A$ or on the predicted joint distribution of states and actions. As a result, our analysis extends without modification to the case in which the agent values information for its own sake (i.e. for non-instrumental reasons).
@@
Formally, let $\Omega$ be the set of $\mathcal{P}(X)$-valued càdlàg functions, let $q:\Omega\times\mathbb{R}_{+}\rightarrow\mathcal{P}(X)$ be the canonical stochastic process on this space, let $\{\mathcal{F}_{t}\}$ be the natural filtration associated with this canonical process, and let $\mathcal{F}=\lim_{t\rightarrow\infty}\mathcal{F}_{t}$. Let $\mathcal{T}\subset\Omega\rightarrow\mathbb{R}_{+}$ be the set of non-negative stopping times with respect to $\{\mathcal{F}_{t}\}$.
@@
$V(P)=\sup_{\tau\in\mathcal{T}}\mathbb{E}^{P}[\hat{u}(q_{\tau})-\bar{\kappa}\tau|\mathcal{F}_{0}].$
@@ 2.2 The Principal’s Problem
The principal chooses the information the agent receives so as to maximize engagement. In the background, we have in mind that the principal is showing ads or otherwise profiting [^7] Network effects, in particular the possibility that this agent’s attention increases utility other agents receive from the platform and thus the profits the principal can extract from those agents, are one relevant engagement maximization motive. from each additional increment of attention the agent pays to the information the principal provides.
@@
We begin with defining a measure for the agent’s information processing using a continuous time version of what Caplin et al. (2022) call a “uniformly posterior-separable” cost function, as described in Hébert and Woodford (2023). Uniformly posterior-separable cost functions are defined in terms of a “generalized entropy function,” $H:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$. We assume that $H$ is strongly convex and twice continuously-differentiable. Using $H$, we define the cumulative information acquired, $I^{H}_{t}$, by
$\frac{\mathrm{d}}{\mathrm{d}t}I_{t}^{H}=\lim\sup_{h\downarrow 0}\frac{1}{h}\mathbb{E}^{P}[H(q_{t+h})-H(q_{t^{-}})|\mathcal{F}_{t^{-}}].$ (1)
We assume the agent can process information at a rate defined by the constraint
$\frac{\mathrm{d}}{\mathrm{d}t}I_{t}^{H}\leq\chi,$ (2)
[…]
@@
The principal’s goal is to maximize “engagement,” which is not necessarily the same as maximizing the information acquired. Specifically, we assume that the principal earns profits in proportion to cumulative information acquisition as measured by the generalized entropy function $G:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$ (another convex and twice continuously-differentiable function). The function $G$ is not necessarily equal to the function $H$, although this will be one case of interest. The former determines the information that the principal would like the agent to acquire; the latter governs the rate at which the agent can acquire different kinds of information. We will assume that there is always some information the principal would like the agent to acquire (i.e. that the Hessian of $G$ is non-zero on the interior of the simplex).
@@
The principal’s goal is to design the agent’s belief process so as to maximize profits, taking into account the fact the agent will know the nature of the beliefs process and optimally choose when to stop paying attention. Let $\bar{q}_{0}\in\mathcal{P}(X)$ be the agent’s prior. The principal chooses his policies from the set $\mathcal{A}(\bar{q}_{0})$, which is the set of probability measures on $(\Omega,\mathcal{F})$ such that $q$ is martingale belief processes with $q_{0}=\bar{q}_{0}$, increasing adapted processes $I^{H}$ with $I_{0}^{H}=0$, and non-negative stopping times $\tau$, such that (1) and (2) are satisfied for all $t\in[0,\tau)$. Formally, the principal chooses the probability measure $P$, which is equivalent to choosing the law of of the belief process $q$. In this sense, the principal can choose any càdlàg martingale belief process with $q_{0}=\bar{q}_{0}$, subject to the constraint imposed by the agent’s information processing capacity. [^9] An alternative interpretation is that the principal can send more information than the agent can process, in which case the agent chooses which information to attend to. Allowing the agent this choice cannot benefit the principal, and it is therefore without loss of generality to suppose the principal chooses a process that satisfies the agent’s information processing constraint.
@@
$J(\bar{q}_{0})$ $=\sup_{(P,I^{H},\tau)\in\mathcal{A}(\bar{q}_{0})}\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})|\mathcal{F}_{0}]$ (P)
@@
$\mathbb{E}^{P}[\hat{u}(q_{\tau})-\bar{\kappa}\tau|\mathcal{F}_{0}]\geq\sup_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\hat{u}(q_{\tau^{\prime}})-\bar{\kappa}\tau^{\prime}|\mathcal{F}_{0}].$

### v6

@@ 1 Introduction
Free-to-use online platforms such as Facebook, Instagram, Youtube, and Pinterest are used by billions of people worldwide and earn substantial profits by displaying advertisements to their users. Their business models are powered by personalized recommendation algorithms that seek to maximize the “engagement” of each user by selectively displaying content (lada_wang_yan_2021, news_feed_engagement). Large quantities of computing and other resources have been invested by these firms to develop algorithms that can predict billions of users’ preferences in real time.
@@
A second, subtler challenge is to manage the algorithm’s own “incentives”. As is suggested by the previous paragraph, to maximize user’s ex ante engagement, platforms often design these algorithms with an initial commitment of providing reliable and wide-ranging content. Subsequent training of the algorithms based on user interactions, however, can induce algorithms to behave in a dynamically inconsistent manner. For instance, at a later time, the algorithm may “learn” to steer users into niche “rabbit holes ” (news_tiktok) because it delivers short-term engagement gains. When the platform lacks long term commitment power, this cause users to question the credibility of the platform’s stated design goals. It is therefore unclear, ex-ante, whether concerns about dynamic inconsistency meaningfully constrain algorithm design.
Engagement maximization is relevant in many other contexts. We will pair the example of an internet platform engaging a user with a second example, that of a teacher who seeks to maximize the engagement of a student who cares only about passing a test. [^1] We are grateful to Emir Kamenica for suggesting this alternative setting as an application of our model. Our theoretical framework applies to this setting without modification.
We study the optimal design of sequential information presentation from a principal-agent perspective. The principal (the platform/teacher) provides the agent (the user/student) with information. The agent derives value from the information, and chooses when to stop engaging with the provided information. [^2] We show that it is without loss of generality to assume the agent does not process the information selectively; that is, the agents attends to the information the principal provides in equilibrium. The rate at which the agent can process information is limited, and the agent experiences an opportunity cost of time spent on the platform net of any utility from using the platform. The principal’s goal is to maximize the attention the agent allocates to the platform (which under some conditions will be equivalent in equilibrium to maximizing the stopping time), and the principal accomplishes this by choosing the nature of the information provided to the agent. The key modeling assumptions we impose are that the principal knows perfectly the agent’s preferences and that the principal can flexibly manipulate the entire information flow.
@@
    Leading Case: All Engagement is Profitable. Suppose that the principal’s engagement measure is identical to the informativeness measure. In the internet platform context, we interpret this case as one in which the agent is exposed to ads in proportion to the context she receives, with no measurement of whether she attends to the ads. In this case, the optimal information structure is narrow but deep: the principal only provides information about the agent’s interested dimensions. This information is provided in a rare but chunky way, leading the agent to spend additional time and acquire additional information relative to the agent-optimal benchmark. This result echoes the observation that engagement maximizing recommendation algorithms often lead users down “rabbit-holes.”
    Leading Case: Decision-Irrelevant Engagement is Profitable. Now suppose that the principal’s revenue comes exclusively from engagement with information that is irrelevant for decision. In the internet platform context, we interpret this case as one in which the agent is exposed to decision-irrelevant ads, and the principal benefits only when the agent engages with these ads (e.g. if ad revenue comes from the click of an ad). In this case, we show that the optimal information structure is broad but noisy: the principal provides excess information (e.g. sponsored content) that the agent views as pure noise, intermixed with just enough decision-relevant information to keep the agent engaged. This result echoes the observation that engagement-maximizing recommendation algorithms often intersperse useful content with “click-bait” and “stealth marketing” material.
@@
    The intuition for the result is simple: We show that the principal-optimal policy leaves the agent with no surplus at the prior. If the agent were to become “more certain than” the prior at an interim stage, the surplus from future information would be negative, and the agent would stop engaging with the principal. On the other hand, if there agent were to become “less certain than” the prior at an interim stage, her surplus from future information would be positive. At this point, the principal would be tempted to adjust the information process and extract more surplus, at the agents’ expense. Thus, the only possible interim beliefs are those that are exactly “as certain as” the prior belief. Our main result in Section 4 is a formalization of this idea, along with a precise definition of what it means to be more certain, less certain, or as certain as the prior.
Our two leading cases are relevant to both the internet platform and teacher-student contexts. In the latter, they map to the degree to which the teacher values the student learning about test-relevant vs. test-irrelevant information. A teacher who values student learning but not test performance will provide just enough test-specific information to engage a student who cares only about the test, while providing additional content that is not part of the test. In contrast, a teacher who values all learning, regardless of whether it is covered on a test, will optimally provide only test-relevant information while inducing students to learn more test-relevant information than they would choose to learn on their own.
@@ 1.1 Related Literature
Our paper contributes to several strands of literature on the dynamic provision of information. Viewing our principal as a media company, our model is related to work on models of media bias (see gentzkow2015media for a survey). We share with kleinberg2022challenge an interest in explaining why the users of internet platforms would engage heavily with those platforms while perceiving themselves as gaining little from doing so. We derive this outcome as a result of strategic behavior by rational agents with conflicting incentives; those authors emphasize the time-inconsistency of user preferences. We share with acemoglu2021misinformation an emphasis on explaining what kind of information is available on internet platforms; our analysis focuses on content selection algorithms, whereas their analysis focuses on information sharing between users.
@@
Closely related to our work is the literature on dynamic Bayesian persuasion (e.g. ely2017beeps; renault2017optimal; ely2020moving; orlov2020persuading; che2020keeping) that build on the static model of kamenica2011bayesian. Our setting differs from most of these papers in that our principal’s payoff solely depends on the engagement of the agent and not on her ultimate choice of action. [^3] The leading case of our model in which all engagement is profitable is equivalent to one where the principal’s only goal is to ensure the agent continues to pay attention, as in kawamura2019news. However, our principal presents unbiased information, and hence is not engaged in cheap talk (crawford1982strategic; cheng2022bayesian). Most of these papers focus on the “Bayesian persuasion” settings where the principal’s payoffs directly depend on the agent’s action and the maximization of “engagement” is a side effect. A more recent strand of literature focuses on the direct maximization of “attention” or “engagement” without the capacity constraint, including two concurrent papers (jan2020dynamic and koh2022attention) and two subsequent papers (koh2024persuasion and saeedi2024getting).444jan2020dynamic’s main focus is on the competition between multiple senders. The single sender case of jan2020dynamic is the closest to ours. koh2022attention and koh2024persuasion focuses on more general implementability characterization. saeedi2024getting focuses on the case with asymmetric prior. These papers share several common predictions with our model, including Poisson dilution signals and the minimization of ex ante user welfare; indicating the robustness of these features across models. However, the absence of an informational constraint in these papers also drives stark differences—these papers predict full revelation of the state upon signal arrival, indicating no bias from the agent’s preferred posterior beliefs. Our analysis emphasizes the beliefs the agent will arrive at, and how these differ between the principal’s optimal strategy and an agent-optimal benchmark, a comparison that is not possible absent information processing constraints. We elaborate further on the connection between our model and these models in Section 5. We assume both players are long-lived and have commitment power, in contrast to the limited commitment settings in orlov2020persuading; che2020keeping or the myopic agent settings in ely2017beeps; renault2017optimal. Interestingly, unlike the findings in these papers, neither commitment power nor forward-lookingness of the agent is necessary for sustaining our equilibrium strategy (see Section 4.2).
@@
Our model predicts gradual information revelation over time, which is a feature shared by many of the dynamic information design models (e.g. ely2015suspense; horner2016selling; che2020keeping; orlov2020persuading). However, unlike these papers, the gradual nature of belief evolution in our model arises from the agent’s information processing constraint, as opposed to a desire to maximize suspense or address problems of limited commitment. In particular, che2020keeping predicts that the agent’s belief involves Poisson jumps and a drift, while our optimal strategy admits Poisson jumps without a drift.
@@
Formally, our approach is a principal-agent version of hebert2019rational. Those authors consider a model in which a single decision maker chooses both what information to acquire and when to stop and act, whereas in our model the principal chooses the information and the agent chooses when to stop and act. We compare our model to a benchmark in which the agent chooses both the information and when to stop and act; this benchmark is characterized by results found in hebert2019rational. We follow hebert2019rational in assuming that the principal can choose any stochastic process for the agent’s beliefs, subject only to the martingale requirement (which is imposed by Bayesian updating) and the upper bound on the agent’s attention. We model this upper bound using a “uniformly posterior-separable” information cost, in the terminology of the rational inattention literature (caplin2017shannon). [^5] Examples of such information costs include mutual information, as applied in sims2010rational, as well as other proposed alternatives (hebert2018informationcosts, bloedel2020cost).
@@
The rest of the paper is organized as follows. We begin in section 2 by describing the basic environment of our model. Section 3 characterizes optimal policy in our baseline model, with an emphasis on the beliefs the agent will hold when stopping. Section 4 discusses the dynamics of beliefs and the role of commitment Section 5 discusses an extension of our baseline model , and in Section 6 we conclude.
@@ 2.1 The Agent’s Problem
We study the problem of a rational, Bayesian agent receiving signals about an underlying state for the purpose of taking an action. We model information acquisition as a continuous time process, building on results in hebert2019rational and zhong2017jumps.
Let $X$ be a finite set of possible states of nature. The state of nature is determined ex-ante, does not change over time, and is not known to the agent. Let $q_{t}\in\mathcal{P}(X)$ denote the agent’s beliefs at time $t\in[0,\infty)$, where $\mathcal{P}(X)$ is the probability simplex defined on $X$. We will represent $q_{t}$ as vector in $\mathbb{R}_{+}^{|X|}$ whose elements sum to one, each of which corresponds to the likelihood of a particular element of $X$, and use the notation $q_{t,x}$ to denote the likelihood under the agent’s beliefs at time $t$ over the true state being $x\in X$. For each state $x$, we use $e_{x}\in\mathcal{P}(X)$ to denote the degenerate belief that state $x$ occurs with probability $1$.
At each time $t$, the agent can either stop and choose an action from a finite set $A$, or continue to acquire information. Let $\tau$ denote the time at which the agent stops and makes a decision, with $\tau=0$ corresponding to making a decision without acquiring any information. The agent receives utility $u_{a,x}$ if she takes action $a$ and the true state of the world is $x$, and pays a flow cost of delay per unit time, $\kappa>0$, until an action is taken. [^6] Note that we assume that the delay cost enters the agent’s utility additively and the agent does not discount the utility from acting. Discounting, in our context, has the effect of changing the relative values of acting and never acting, and for this reason would complicate our analysis; see the more extensive discussions in hebert2019rational and zhong2017jumps on this point.
Let $\widehat{u}(q^{\prime})$ be the payoff (not including the cost of delay) of taking an optimal action under beliefs $q^{\prime}\in\mathcal{P}(X)$:
$\widehat{u}(q^{\prime})=\max_{a\in A}\sum_{x\in X}q^{\prime}_{x}u_{a,x}.$
In what follows, the convex function $\widehat{u}:\mathcal{P}(X)\rightarrow\mathbb{R}$ will summarize the value the agent places on information. This function, as opposed to the utility function $u_{a,x}$, can be viewed as the primitive “value of information when acting” in our model; nothing in our analysis will depend on the nature of the action space $A$ or on the predicted joint distribution of states and actions. As a result, our analysis extends without modification to the case in which the agent values information for its own sake (i.e. for non-instrumental reasons).
@@
Formally, let $\Omega$ be the sample space. Let $q:\Omega\times\mathbb{R}_{+}\rightarrow\mathcal{P}(X)$ be a canonical càdlàg stochastic process on $\mathcal{P}(X)$, let $\{\mathcal{F}_{t}\}$ be the natural filtration associated with this canonical process, and let $\mathcal{F}=\lim_{t\rightarrow\infty}\mathcal{F}_{t}$. Let $\mathcal{T}$ be the set of non-negative stopping times with respect to $\{\mathcal{F}_{t}\}$.
@@
$V(P)=\sup_{\tau\in\mathcal{T}}\mathbb{E}^{P}[\widehat{u}(q_{\tau})-\kappa\tau|\mathcal{F}_{0}].$
@@ 2.2 The Principal’s Problem
The principal chooses the information the agent receives so as to maximize engagement. We begin with defining a measure for the agent’s information processing using a continuous time version of what caplin2017shannon call a “uniformly posterior-separable” cost function, as described in hebert2019rational. Uniformly posterior-separable cost functions are defined in terms of a “generalized entropy function,” $H:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$. We assume that $H$ is twice continuously-differentiable and strongly convex. [^7] I.e. the Hessian matrix of $H$ is strictly positive definite. We assume the agent can process information at a rate no greater than $\chi>0$. That is, the process $H(q_{t})-\chi t$ must be an $\mathcal{F}_{t}$-supermartingale, requiring that for all $t$ and $\delta>0$,
$\mathbb{E}^{P}[H(q_{t+\delta})-\chi(t+\delta)|\mathcal{F}_{t}]\leq H(q_{t})-\chi t,$ (1)
where we adopt the convention that $q_{t}=q_{\tau}$ for $t>\tau$. [^8] This constraint is equivalent to the one studied in hebert2019rational within the class of uniformly posterior-separable costs, allowing us to invoke their results.
@@
The principal’s goal is to maximize “engagement,” which is not necessarily the same as maximizing the information acquired. Specifically, we assume that the principal earns profits in proportion to cumulative information acquisition as measured by the convex function $G:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$. The function $G$ is not necessarily strictly or strongly convex, although $G=H$ will be one of our leading cases.
@@
The principal’s goal is to design the agent’s belief process so as to maximize profits, taking into account the fact the agent will know the nature of the beliefs process and optimally choose when to stop paying attention. Let $\bar{q}_{0}\in\mathcal{P}(X)$ be the agent’s prior. The principal chooses his policies from the set $\mathcal{A}(\bar{q}_{0})$, which is the set of probability measures on $(\Omega,\mathcal{F})$ such that $q$ is martingale belief processes with $q_{0}=\bar{q}_{0}$ and non-negative stopping times $\tau$ such that (1) is satisfied. Formally, the principal chooses the probability measure $P$, which is equivalent to choosing the law of the belief process $q$. In this sense, the principal can choose any càdlàg martingale belief process with $q_{0}=\bar{q}_{0}$, subject to the constraint imposed by the agent’s information processing capacity. [^9] An alternative interpretation is that the principal can send more information than the agent can process, in which case the agent chooses which information to attend to. Allowing the agent this choice cannot benefit the principal, and it is therefore without loss of generality to suppose the principal chooses a process that satisfies the agent’s information processing constraint.
@@
$J(\bar{q}_{0})$ $=\sup_{(P,\tau)\in\mathcal{A}(\bar{q}_{0})}\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})|\mathcal{F}_{0}]$ (P)
@@
$\mathbb{E}^{P}[\widehat{u}(q_{\tau})-\kappa\tau|\mathcal{F}_{0}]\geq\sup_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\widehat{u}(q_{\tau^{\prime}})-\kappa\tau^{\prime}|\mathcal{F}_{0}].$